NDC 24208-750-06 FOR OPHTHALMIC USE ONLY KEEP OUT OF REACH OF CHILDREN Each ml contains Active: Atropine Sulfate USP, 1% (10If the nurses administer the maximum number of doses of atropine per day, for how many whole days will the bottle of atropineIII. What is the ratio strength concentration of benzalkonium chloride in the atropine ophthalmic solution?
Answer #1

1.

1 ml =10 mg atropine= 25 drops

1drop= 10/25=0.4 mg

Ordered dose= 0. 4mg every 4 hourly orally

ie, 1drop 6 times a day

ie, total 6 drops per day orally

2.

Maximum drops to administer= 6 drops

Total volume of solution= 15 ml

1ml= 25 drops

15 ml= 25*15= 375 drops

One day 6 drops is maximum

So 375 drops can be administered in = 375/6= 62.5= 62 complete days.

3.

1%=10mg

0.01%= x

1/10= 0.01/x

X= 0.01*10= 0.1 mg

So, 1ml contain 0.1 mg benzalkonium chloride